# Lightweight Neural App Control

### v1

@@ 6 Conclusion
One of the main limitations of the proposed method is the limited training data. LiMAC is trained on just 13K and 18K episodes for AndroidControl and AitW, respectively. The absence of any pretraining further hinders the model’s ability to improve performance on more complex tasks. In the future, we aim to enhance the model’s performance by incorporating online learning techniques, such as reinforcement learning. After the initial training stage presented in this work, LiMAC could interact with an Android emulator to generate additional data. By using a suitable reward function, or even leveraging GPT-4 to evaluate the generated trajectories and assign rewards (Bai et al., 2024), we could fine-tune LiMAC to improve the completion rate of tasks.
@@ Appendix A Dataset Format
Observation: The exact form of the observation depends on the type of model it is used for. Text-based approaches such as T3A need a textual observation. For AndroidControl, we use the provided accessibility UI trees, which we further process into a list of UI elements containing information such as the element type, description, and attributes (clickable, editable, selected, etc…). Like in Li et al. (2024), we filter these to retain only important elements, namely those that contain text or have critical attributes. For AitW, OCR representations of text and icons are given in the dataset, but no comprehensive UI trees are provided. Therefore, to obtain the final representation, each element must be identified and converted using a similar procedure to that for AndroidControl. Vision models such as Qwen2-VL and Florence2 will expect an image-based observation. This observation will consist of the current phone screenshot along with an overlay of the UI element bounding boxes and their index. Finally, some models, such as M3A and ours, use a mixture of observations, both text and image-based. In particular, our model expects a text-based list of UI elements similar to the one described above, as well as a list of cropped images. The list of cropped images corresponds to each of the UI elements in the text-based observation and is used by our model as described in Section 3.1.
Action: Action grounding is a crucial part of mobile phone control, so as in previous works (Zheng et al., 2024; Yang et al., 2023; Wang et al., 2024b; a) and both the datasets we use, we define a fixed action space, seen in Table 5. Information for most actions is sourced directly from the datasets, with only the action name at times varying. The only exceptions to this are the click and long-press actions, which require a target element, rather than x-y coordinates. For these, we select the best matching candidate from the observation list of UI elements. The action takes a specific JSON format we expect the models to match to facilitate parsing, which is simply a dictionary with the action type and an action specification (see Table 5). An example would be: {"action-type":"open-app","app-name":"Chrome"}.
Table 5: Agent action space. open-app, wait, and long-press do not feature in AitW, while navigate-home does not feature in AndroidControl.
Action type Action specification
@@ Appendix C Implementation Details
Appendix D Additional Results
This section presents additional evaluation results for LiMAC. In Table 6, we provide a full set of evaluation metrics for the baseline models, as well as for various combinations of LiMAC with other methods. These combinations are used to predict the target element in click actions or generate text for specific actions, such as open-app and input-text. In all the experiments involving LiMAC, AcT is employed to predict the action type, while different combinations of methods are used to predict the action specifications, such as the target element or text generation. This approach allows us to isolate the impact of each combination on performance while maintaining a consistent action type prediction. This table extends the results already presented in Tables 1, 2 and 3 providing a more in-depth understanding of the performance across a range of metrics. This additional breakdown offers a clearer understanding of how LiMAC performs when integrated with other methods, offering insights into the strengths and potential trade-offs of each combination in different scenarios.
@@ Appendix E Case Studies
Some sample episodes from AndroidControl, including agent predictions, are provided in Figures 4 and 5. These are provided for illustration purposes, as well as to further explain ‘relaxed’ accuracies and an example failure. Figure 4 presents both an instance of a relaxed target element in the third timestep and a failed input-text action in the final timestep. Figure 5 shows a relaxed input-text action in the fourth timestep and an otherwise successful episode. Further details are provided in the figure captions.

### v2

@@ 6 Conclusion
One of the main limitations of the proposed method is the limited training data. LiMAC is trained on just 13K and 18K episodes for AndroidControl and AitW, respectively. The absence of any pretraining further hinders the model’s ability to improve performance on more complex tasks. In the future, we aim to enhance the model’s performance by incorporating online learning techniques, such as reinforcement learning. After the initial training stage presented in this work, LiMAC could interact with an Android emulator to generate additional data. By using a suitable reward function, or even leveraging GPT-4 to evaluate the generated trajectories and assign rewards (Bai et al., 2024), we could fine-tune LiMAC to improve the completion rate of tasks. An important focus for future work will be to develop error handling and recovery mechanisms to enable high success rates and robustness in online interactions. Another area of future research could address the safety of such models when handling sensitive data, such as credit card information and personal identifiers. It is essential to design foundation models with robust security protocols to protect against data breaches, especially when interacting with mobile phones containing sensitive information.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant Nos. 62422605, 92370132).
@@ Appendix A Dataset Format
Observation: The exact form of the observation depends on the type of model it is used for. Text-based approaches such as T3A need a textual observation. For AndroidControl, we use the provided accessibility UI trees, which we further process into a list of UI elements containing information such as the element type, description, and attributes (clickable, editable, selected, etc…). Like in Li et al. (2024a), we filter these to retain only important elements, namely those that contain text or have critical attributes. For AitW, OCR representations of text and icons are given in the dataset, but no comprehensive UI trees are provided. Therefore, to obtain the final representation, each element must be identified and converted using a similar procedure to that for AndroidControl. Vision models such as Qwen2-VL and Florence2 will expect an image-based observation. This observation will consist of the current phone screenshot along with an overlay of the UI element bounding boxes and their index. Finally, some models, such as M3A and ours, use a mixture of observations, both text and image-based. In particular, our model expects a text-based list of UI elements similar to the one described above, as well as a list of cropped images. The list of cropped images corresponds to each of the UI elements in the text-based observation and is used by our model as described in Section 3.1.
Action: Action grounding is a crucial part of mobile phone control, so as in previous works (Zheng et al., 2024; Yang et al., 2023; Wang et al., 2024b; a) and both the datasets we use, we define a fixed action space, seen in Table 5. Of this action space, open-app, wait, and long-press do not feature in AitW, while navigate-home does not feature in AndroidControl. Information for most actions is sourced directly from the datasets, with only the action name at times varying. The only exceptions to this are the click and long-press actions, which require a target element, rather than x-y coordinates. For these, we select the best matching candidate from the observation list of UI elements. The action takes a specific JSON format we expect the models to match to facilitate parsing, which is simply a dictionary with the action type and an action specification (see Table 5). An example would be: {"action-type":"open-app","app-name":"Chrome"}.
Table 5: Agent action space, along with the relevant datasets.
Action type Action specification AitW AndCtrl
@@ Appendix C Implementation Details
Appendix D Additional Studies
This section presents additional evaluation results for LiMAC.
D.1 Extended Success Rate Table
In Table 6, we provide a full set of evaluation metrics for the baseline models, as well as for various combinations of LiMAC with other methods. These combinations are used to predict the target element in click actions or generate text for specific actions, such as open-app and input-text. In all the experiments involving LiMAC, AcT is employed to predict the action type, while different combinations of methods are used to predict the action specifications, such as the target element or text generation. This approach allows us to isolate the impact of each combination on performance while maintaining a consistent action type prediction. This table extends the results already presented in Tables 1, 2 and 3 providing a more in-depth understanding of the performance across a range of metrics. This additional breakdown offers a clearer understanding of how LiMAC performs when integrated with other methods, offering insights into the strengths and potential trade-offs of each combination in different scenarios.
@@ D.2 Confusion Matrix
D.3 Failure Analysis
We also examine the failure patterns of LiMAC, using Florence2 as the VLM, across the two datasets studied. Figure 4 displays the frequency of these failures, categorised by the type of failure in predicting either the action type or the action specifications. Specifically, within the action specifications, failures occur in two areas: incorrect prediction of the click target and inaccurate generation of input text by the VLM. In both datasets, the most common type of failure is misclassification of the action type, closely followed by failures in predicting the click target. These findings underscore the key challenges that research on app control should address.
Figure 4: Relative frequency of different types of action prediction errors in the two datasets
D.4 UI Elements Scalability
In this section, we assess how the number of UI elements in an observation impacts the success and failure rates of action prediction. Figure 5 displays the number of successful and unsuccessful action predictions made by LiMAC, categorised by the number of UI elements. The results are grouped into bins of ten on the x-axis. The number of UI elements extends up to 150 in AitW and up to 290 in AndroidControl. However, for clarity, only bins containing more than five samples are included in the figures. Overall, the data suggests that the rate of failed action predictions increases slightly as the number of UI elements grows. This trend is expected since accurately predicting the target of click actions becomes more challenging with more UI elements present.
Figure 5: Number of successful and failed prediction of actions with respect to the number of UI elements in the observation, for the two datasets.
@@ Appendix E Case Studies
Some sample episodes from AndroidControl, including agent predictions, are provided in Figures 6 and 7. These are provided for illustration purposes, as well as to further explain ‘relaxed’ accuracies and an example failure. Figure 6 presents both an instance of a relaxed target element in the third timestep and a failed input-text action in the final timestep. Figure 7 shows a relaxed input-text action in the fourth timestep and an otherwise successful episode. Further details are provided in the figure captions.